a welding fuel gas contains carbon and hydrogen only. burning a small sample of it in oxygen gives 3.38 g carbon dioxide, 0.69 g of water and no other products. A volume of 10L(measured at STP) of this welding gas is found to weigh 11.6 g. calculate its empirical formula

Grade:12

1 Answers

Vikas TU
7 years ago
moles at STP = 10/22.4 => 0.44642
Apply POAC on carbon after,
 
0.44642 * x = 1 * 3.38/44
x = 0.076818/0.44642
  = 0.172076
for hydrogn u will get too 0.172076
Hence EPIROCAL formulae becomes=>
0.1720(CH)
=> CH.